monochromatic light of wavelentgh 589nm isincident from air to to a water surface. if refractive index is 1.33 find the wavelength, frequency & speed of light
 

Grade:12

1 Answers

Arun
5 years ago

Wavelength of incident monochromatic light,

λ = 589 nm = 589 × 10^−9 m

Speed of light in air, c = 3 × 10^8 m/s

Refractive index of water, μ = 1.33

(a) The ray will reflect back in the same medium as that of incident ray. Hence, the wavelength, speed, and frequency of the reflected ray will be the same as that of the incident ray.

Frequency of light is given by the relation,

v = c / λ

   = 3 x 10^8 / 589 x 10^-9

   = 5 .09 x 10^14 Hz

Hence, the speed, frequency, and wavelength of the reflected light are 3 × 10^8 m/s, 5.09 ×10^14 Hz, and 589 nm respectively.

 

 

 

(b) Frequency of light does not depend on the property of the medium in which it is travelling. Hence, the frequency of the refracted ray in water will be equal to the frequency of the incident or reflected light in air.

So Refracted frequency, ν = 5.09 ×10^14 Hz

Speed of light in water is related to the refractive index of water as:

v = c / μ

v =  3 x 10^8 / 1.33 = 2.26 x 10^8 m/s 

 

Wavelength of light in water is given by the relation,

 λ = v / v

v =  2.26 x 10^8 / 5 .09 x 10^14 = 444.007 x 10^-9 m

   = 444.01 nm

Hence, the speed, frequency, and wavelength of refracted light are 2.26 ×10^8 m/s, 444.01nm, and 5.09 × 10^14 Hz respectively